OpenStudy (anonymous):

a train accelerates from rest (0 m/s) at a rate of 3km/hr per second. What is the train's speed after 12 seconds?

OpenStudy (anonymous):

use the formula,v=u+ft

OpenStudy (anonymous):

here,v=the speed after 12 seconds..t=12, u=0 and f=3000/3600 m/second

OpenStudy (anonymous):

Here, initial speed is 0 so : u = 0.. So, using v = u + at a = 3*(5/18) = 5/6 and t = 12 : v = 0 + (5/6)*12 v = 5*2 = 10 m/s

OpenStudy (anonymous):

\(\color{blue}{v = 10 \; m/s}\)